OpenStudy (anonymous):

What is the slope of the line that passes through (2, 12) and (4, 20)

OpenStudy (saifoo.khan):

\[m = \frac{20-12}{4-2}\] \[m = \frac{8}{2}\] m = ????

OpenStudy (anonymous):

(20-12)/(4-2) =